A box have 20 tickets, including 4 winning tickets. Two people take 2 tickets from the box randomly. calculate the probability that each person will get at least 1 winning ticket (using counting, permutations and combinations)
Answer #1

Below 3 Scenarios are possible:

2W and 2 L

3 W and 1 L

4 W

P(1) = (4C2 * 16C2 * ½)/20C4 = 0.074

P(2) = 4C3 /20C4 = 0.001

P(3) = 4C4 /20C4 = 0.0002

Hence, P = P(1) + P(2) + P(3) = 0.075 or 7.5%